SLAC PEP-II
BABAR
SLAC<->RAL
Babar logo
HEPIC E,S & H Databases PDG HEP preprints
Organization Detector Computing Physics Documentation
Personnel Glossary Sitemap Search Hypernews
Unwrap page!
Det. Search
Who's who?
Meetings
FAQ
Images
Archive
Systems
Performance
Intern. region
Vertex Tracker
Drift chamber
DIRC
Calorimeter
IFR
LST
Magnet
Electronics
Trigger
Operations
Run Coordination
Contact Experts
Shift Takers Info
Operations Manual
Electronic Logbook
Ops Hypernews
Shift Signup
Check this page for HTML 4.01 Transitional compliance with the
W3C Validator
(More checks...)

Minutes of the Trigger Operation Meeting (21st August, 2007)

Attendees: Debbie Bard, Rainer Bartoldus, Al Eisner, Kim Hojeong, Rahmat Rahmats.

Operations

  • Sunday swing shift: email for GLT timing plot bin 1 (IFT) low by ~1tick. Did not appear so bad in L1 LiveFastMon file so possibly an artifact of low statistics. Disappeared in next run.
  • Paged Monday AM for ODF damage - MPIC spurious interrupt error in ROM log file so instructed pilot to power-cycle crate. This is the first time we've seen this error in a while - frequency seems to be falling off after a busy July.
  • Please take a look at the slides for the ~monthly TRG status report for tomorrow's 3.45 meeting.

EMT

  • Jim's question about the EMT calibration remains - should look at code to check if it always returns "success".

L3

  • Validation of Arik's code changes are in progress.
    • In data: looking for losses in single/double track lines. Also checking lines used to calculate hadronic ratio so that lumi ratio does not change when the code is introduced.
      -> Done! changes small.
    • In MC: Ensure we can adequately simulate old data- the old track-finding should not change with the code modifications alone!
  • Can turn code on this month if we know lumi is not affected.

Data Quality and Monitoring

  • All is well with the data so far. Other work is still in progress.

Trigger Configuration

  • Study is still on-going. See status page for up-to-date information.
  • The 1E line could be put in this week. Still need to check whether it will affect other open lines - e.g. Bhabha rates could blow up.

This page is created by Debbie Bard (djbard@slac.stanford.edu); however, you are welcome to edit and/or correct above items, specially for those who attended the meeting. Thanks.